The greater of two consecutive integers is 10 ess than twice the smaller. Find the integers.
(I'm guessing 2(2x+1)-10=2x+5, but it doesn't work out...)

Answer by checkley77(12844)

2(2x+1)-10=2x+5
WHERE DO YOU GET THE 2(2X+1) & THE +5?
------------------------------------------
LET X & (X+1) BE THE 2 INTEGERS.
X+1-10=2X
X-2X=10-1
-X=9
X=-9 ANS. FOR THE SMALLER INTEGER.
-9+1=-8 FOR THE LARGER INTEGER.
-8-10=2*-9
-18=-18